# Algebraic Independence of Special Points on Shimura Varieties

### v1

@@ 3. Bounds on the Galois orbits
Fix a choice of fundamental domain $\mathcal{F}_{S}$ for the uniformization $q:X\to S$ of the Shimura variety $S$. Let $F$ be the number field over which $S$ admits a canonical model. The degree of $F$ is bounded in terms of the datum $(G,X)$ and $K$. All the special points of $S$ are algebraic points defined over abelian extensions of $F$.
For a special point $s\in S$, let $x$ be a preimage of $s$, i.e. $x$ is a pre-special point. By definition, the Mumford-Tate group of a special point is an algebraic torus. Let $K^{m}_{\mathbf{M}}$ be the maximal compact open subgroup of $\mathbf{M}(\mathbb{A}_{f})$ and $K_{\mathbf{M}}$ the compact open subgroup $K\cap\mathbf{M}(\mathbb{A}_{f})$ of $\mathbf{M}(\mathbb{A}_{f})$. Let $E$ be the splitting field of $\mathbf{M}$. Since $G$ is of adjoint type, $E$ is a (Galois) CM field. Let $D_{E}$ be the absolute value of the discriminant of $E$.
@@
Define the discriminant of $s$ as
$\Delta(s)=[K^{m}_{\mathbf{M}}:K_{\mathbf{M}}]D_{E}.$
We want to estimate the heights and degrees of special and pre-special points in terms of $\Delta(s).$ Fortunately, recent research progress allows us to have the following proposition.
@@
Where the constants $C_{1},C_{2},\cdots$ are depend on $G,X,F,\mathcal{F}_{S}$ and the realization.
@@
Suppose $T$ is a semi-abelian variety over a number field $L$, and $K$ is any number field. Let $q$ denote the Néron-Tate height on $T$ defined with respect to some symmetric ample divisor and let
$\eta:=\eta(T,L)=\operatorname{inf}q(P),$
where the infimum is taken over all non-torsion $P$ in $T(K)$, and we write
$\omega:=\omega(T,K)$
for the cardinality of the torsion group of $T(K)$. Note that $\eta$ and $\omega$ depend on the embedding of $T$ in a projective space. Suppressing this dependence in our notation, we have the following theorem of Masser [Mas88, Theorem $\mathbb{G}_{m},A$].
@@
Suppose $P_{1},\ldots,P_{n}$ on $T(K)$ have Néron-Tate heights at most $q\geq\eta$. Then the relation group of $P_{1},\ldots,P_{n}$ generated by
$\boldsymbol{m}:m_{1}P_{1}+\cdots m_{n}P_{n}=0_{T}$
@@
General semi-abelian subvarieties of $T$ are generated by relation groups where the $m_{i}\in\operatorname{End}(T)$. We accommodate for $\operatorname{End}(T)\neq\mathbb{Z}$ in the following way. Let $\{1,\tau_{1},\dots,\tau_{k}\}\in\operatorname{End}(T)$ be a set of generators. Then, we can work in $T^{(k+1)n}$ with $P_{i},\tau_{1}P_{i},\cdots\tau_{k}P_{i}$ for $1\leq i\leq n$. Due to the dominance of a height associated to an ample divisor (see [Lan83, Chap. 4, Prop. 5.4] or [Poo01, Lemma 6]),
(3.0.1) $q(\tau_{i}(P))\leq C^{i}q(P)$
where each $C^{i}$ is an absolute constant depending on $\tau_{i}$ and the embedding of $T$. Let
@@
where $\overline{C}$ is taken as the maximal of $(C^{i})^{(2ng-1)/2}$ and $q:=\textup{max}_{1\leq i\leq n}\{q(P_{i})\}$.
To estimate 3.0.2 in terms of degree of the algebraic points, we need some estimates for $\eta,\omega$, which is given in [Mas88]. Let $D=[K:\mathbb{Q}]$ and $L=\log(D+2)$ we have:
@@
As to bound the height of rational points on a suitable definable set in Section 4, we need the following upper bound of the norm of the generating set for the linear relations satisfied by the $n$ fixed $V$-images of special points in terms of their discriminants.
@@
Let $(s_{1},x_{1}),\cdots,(s_{n},x_{n})\in S\times T\times\mathbb{G}_{a}^{m}$ be $V$-graphs of special points with discriminants $\Delta$ and define the complexity of $s=(s_{1},\cdots,s_{n})$ by
@@
Then there are constants $C,C^{\prime},c^{\prime}$ depending on $S,T,V,n$ such that for $\boldsymbol{\Delta}\geq C$, there is a generating set for the linear relations satisfied by the $x_{i}\in T$ such that
@@
Since $V$ is a correspondence in $S\times T$, $(s,x)\in V$ is an algebraic point, using standard properties of heights we have that
@@
and since the projections of $V$ are finite dominant to each factors, we have
@@
It is a standard fact that provided the embedding corresponds to a symmetric divisor on $T$, the differences of $q(P)$ and $h(P)$ are bounded by an absolute constant $\delta$ that does not depend on $L^{\prime}$ whiles $P$ goes through $T(L^{\prime})$. Therefore for $h:=\textup{max}\{h(x_{1}),\cdots,h(x_{n})\}$ sufficiently large we have
@@
By Proposition 3.2(a) $\eta,q$ is bounded in terms of $\boldsymbol{\Delta}$, and the proposition follows from Corollary 3.4.
@@ 4.1. Proof of the main theorem over $\overline{\mathbb{Q}}$
First, we prove two simple lemmas.
@@
Let $d=\dim W$. We can find a generic point $P_{1}\times P_{2}\times\cdots\times P_{m+1}\in W^{m+1}$ such that $\operatorname{tr.deg.}_{\mathbb{Q}}\mathbb{Q}(P_{1},\dots,P_{m+1})=d(m+1)$, and we claim that these $P_{i}$ satisfy the above property. The forward direction of the if and only if is clear. Now suppose that $P_{i}\times Q\in V$ for all $i$. Viewing $P_{i}$ as a set, let $\tilde{P_{i}}\subset P_{i}$ be a minimal transcendental basis for $\operatorname{tr.deg.}_{\mathbb{Q}}\mathbb{Q}(P_{i})$. Write $Q=(q_{1},\dots,q_{n})\in\mathbb{C}^{n}$ and let $Q_{i}\subset\{1,2,\dots,n\}$ be the set of indices $j$ so that $q_{i}$ is algebraic over $\mathbb{Q}(\tilde{P_{i}})$. By construction, $\lvert\tilde{P_{i}}\rvert=d$ and the $\tilde{P_{i}}$ are algebraically independent. This implies that $Q_{i}\cap Q_{j}=\varnothing$ for $i\neq j$ and so $\sum_{i=1}^{m+1}\lvert Q_{i}\rvert\leq m$. Thus, there exists an $i$ such that $\lvert Q_{i}\rvert=0$ and for this $i$ we have
@@
Let $S=\overline{\{P_{i}\}}^{\mathrm{Zar}}$ and $T=\overline{\{Q\}}^{\mathrm{Zar}}$ be the $\overline{\mathbb{Q}}$-Zariski closure of $P_{i}$ and $Q$. Then $\dim S=\operatorname{tr.deg.}_{\mathbb{Q}}\mathbb{Q}(P_{i})$ and $\dim Q=\operatorname{tr.deg.}_{\mathbb{Q}}\mathbb{Q}(Q)$. Note that the Zariski closure of $P_{i}\times Q$ has dimension $\operatorname{tr.deg.}_{\mathbb{Q}}\mathbb{Q}(P_{i},Q)=\dim S+\dim T$, showing that the $\overline{\mathbb{Q}}$-Zariski closure of $P_{i}\times Q$ is $S\times T$. But $S\times T\subset V$ by definition and $S=W$ showing inclusion in the other direction. ∎
@@
Suppose $V\subset S\times T$ is an algebraic subvariety of a Shimura variety. There exists a finite set $\Sigma$ of sub-Shimura datum $(H,X_{H})$ of $(G,X)$ and splittings $(H^{\operatorname{ad}},X_{H}^{\operatorname{ad}})=(H_{1},X_{1})\times(H_{2},X_{2})$ such that if $W\subset V$ is exemplary, then there exist one such splitting in $\Sigma$ such that $\pi_{S}(W)$ is the image of $X_{1}\times\{x_{2}\}$, for some $x_{2}\in X_{2}$, under the uniformization map of $S$.
Moreover, if $\mathcal{V}\to W$ is a definable family of $V\subset S\times T$, then the result still holds.
@@
The result for a fixed $V\subset S\times T$ is proven by [PT22, Prop. 3.4] and [DR18, Prop. 6.10]. The version for family is given by the uniform Ax–Schanuel theorem, which is implied by the Ax–Schanuel theorem. This is proven in [ES23, Prop. 2.20]. ∎
@@
Let $S$ be a connected Shimura variety, $T$ be a commutative connected algebraic group. Suppose $V\subset S\times T$ is a proper irreducible subvariety that is finite over $S$. Moreover, suppose that $T$ and $V$ are defined over $\overline{\mathbb{Q}}$. Then, there are only finitely many exemplary components in $V$.
@@
Let $\pi_{S}\colon V\to S$ and $\pi_{T}\colon V\to T$ be the projections onto the two factors. Let $W\subset V$ be an exemplary component and let $S^{\prime}\coloneqq\pi_{S}(W)$ be the special subvariety of $S$ which $W$ maps onto, and let $T^{\prime}\supset\pi_{T}(W)$ be the smallest special subvariety of $T$ containing $\pi_{T}(W)$. $S^{\prime}$ is a special subvariety of $S$ and hence there is a sub-Shimura datum $(H,X_{H})$ of $(G,X)$, a decomposition $(H^{\operatorname{ad}},X_{H}^{\operatorname{ad}})=(H_{1},X_{1})\times(H_{2},X_{2})$, and a point $y_{2}\in X_{2}$ so that $S^{\prime}$ is the image of $X_{1}\times\{y_{2}\}$. By Lemma 4.2, there are only finitely many choices of $X_{H}$ and $X_{1}$. Thus, to show finiteness of exemplary components, it suffices to show that for a fixed $X_{1}$ and fixed dimension of $T^{\prime}$, there are only finitely many points $y_{2}\in X_{2}$ arising from exemplary subvarieties.
Let $\tilde{q}\colon X_{1}\times X_{2}\to X\to S$ denote the uniformization map of $S$ restricted to $X_{1}\times X_{2}$ and let $q\colon F_{1}\times F_{2}\to S$ denote the restriction of $\tilde{q}$ to a fundamental domain $F_{1}\times F_{2}$ of $X_{1}\times X_{2}$. Let $\exp\colon\mathbb{C}^{g}\to T$ denote the uniformization map of the algebraic group and let $e\colon F_{T}\to T$ denote the restriction of that map to a fundamental domain. Using Lemma 4.1, let $x_{1},\dots,x_{r}\in F_{1}$ denote a set of points to determine if $q(X_{1}\times\{x_{2}\})\times\exp(z)\subset V$. Let $\xi_{1},\dots,\xi_{g}$ denote the coordinates in $T$ and let $\tau_{1},\dots,\tau_{k}\in\operatorname{End}(T)$ be a $\mathbb{Z}$-basis for $\operatorname{End}(T)$. Take a generating set of all equations of the form
@@
that all the points of $T^{\prime}$ satisfy, and let $T_{0}$ be the identity component of the algebraic subgroup of $T$ defined by these equations. We can extend each map $\tau_{i}\colon T\to T$ to an endomorphism of its covering space $\tilde{\tau_{i}}\colon\mathbb{C}^{g}\to\mathbb{C}^{g}$ satisfying $\tilde{\tau_{i}}(0)=0$. Suppose that $T_{0}$ is cut out by $\ell$ such equations. Let
$Y=$$\left\{(y,z_{1},\dots,z_{g},m_{111},\dots,m_{gk\ell},b_{1},\dots,b_{\ell})\in F_{2}\times F_{T}\times\mathbb{R}^{gk\ell}\times\mathbb{R}^{\ell}:\right.$
$\left.\forall 1\leq i\leq r,(q(x_{i},y),\exp(\boldsymbol{z}))\in V\text{ and }\forall 1\leq l\leq\ell,\sum_{i,j}m_{ijl}\tilde{\tau_{j}}(z_{i})=b_{l}\right\}$
@@
and set $Z$ to be the projection of $Y$ to $F_{2}\times\mathbb{R}^{gk\ell}\times R^{\ell}$. Both $Y$ and $Z$ are definable sets. The set $Y$ parametrizes points $y\in F_{2}$ so that a $V$-image of $X_{1}\times\{y\}$ lies within a proper special subvariety cut out by the $\boldsymbol{m}$ of $T$ (but not all choices of $\boldsymbol{m}$ and $\boldsymbol{b}$ correspond to algebraic subvarieties).
@@
Suppose for the sake of contradiction that there were infinitely many $\overline{\mathbb{Q}}$-algebraic exemplary subvarieties $W^{\prime}$ with fibers over $X_{1}\times X_{2}$, and with dimension of the smallest special subvariety of $T$ containing $\pi_{T}(W^{\prime})$ equal to a fixed dimension $d=\dim T^{\prime}$. Each one gives a $\overline{\mathbb{Q}}$-point $(y^{\prime},\boldsymbol{m},\boldsymbol{b})\in Z$, and the $\operatorname{Gal}(\overline{\mathbb{Q}}/K)$-orbits of $W^{\prime}$ also lie in $Z$ for $K$ the defining field of $S,T,\operatorname{End}(T)$. Over $\overline{\mathbb{Q}}$, any commutative connected group $T$ can be written as a product of a semi-abelian variety with affine space $B\times\mathbb{G}_{a}^{n}$ (see [NW14, Prop. 5.1.12]). The tuple $\boldsymbol{m}$ consists integers and for semi-abelian varieties, we may take the fundamental domain so that the real part of each $z_{i}$ is in the interval $[0,1]$. Thus, each $b_{l}$ corresponding to an equation on the semi-abelian variety is an integer bounded by $\sum\lvert m_{ijl}\rvert$, which by Proposition 3.5, is bounded by $C\lvert\Delta(y)\rvert^{c}$. Special subvarieties of $\mathbb{G}_{a}^{n}$ are given by the equations $\sum m_{ij}\tilde{\tau_{j}}(z_{i})=0$, so $b_{l}=0$ for those equations. Thus, taking the Galois orbit of $W^{\prime}$ gives a point with different $y^{\prime}$ as well as different $b_{l}$, but the same $\boldsymbol{m}$. View $Z$ is fibered over the $\boldsymbol{m}$ variable and let $\Sigma\subset Z_{\boldsymbol{m}}$ be the set of points arising from exemplary subvarieties $W^{\prime}$ as well as their Galois conjugates. By Proposition 3.2, there are at least $C^{\prime}\lvert\Delta(y^{\prime})\rvert^{c^{\prime}}$ points of height less than $C\lvert\Delta(y^{\prime})\rvert^{c}$. Then Theorem 4.3 gives the existence of a set $R\subset Z_{\boldsymbol{m}}$ of positive dimension whose projection to $F_{2}$ is connected semialgebraic and whose projection to $\mathbb{R}^{\ell}$ is non-constant.
@@
Let $\overline{\exp}\colon\mathbb{C}^{g}/\tilde{T_{0}}\to T/T_{0}$ be the exponentiation map of $T/T_{0}$. Let $F_{T_{0}}$ denote the image of the fundamental domain $F_{T}$ under the quotient map, which will serve as a fundamental domain for $\overline{\exp}$. First, take the preimage of $R\subset Z_{\boldsymbol{m}}$ under the projection map $Y\to Z$ and then project the preimage under the map
$Y_{\boldsymbol{m}}\to F_{2}\times F_{T_{0}}\times\mathbb{R}^{\ell}\to F_{2}\times F_{T_{0}}.$

### v2

@@ 3. Bounds on the Galois orbits
Fix a choice of fundamental domain $\mathcal{F}_{S}$ for the uniformization $q:X\to S$ of the Shimura variety $S$. Let $F$ be the number field over which $S$ admits a canonical model. The degree of $F$ is bounded in terms of the datum $(G_{S},X)$ and $K$. All the special points of $S$ are algebraic points defined over abelian extensions of $F$.
For a special point $s\in S$, let $x$ be a preimage of $s$, i.e. $x$ is a pre-special point. By definition, the Mumford–Tate group of a special point is an algebraic torus. Let $K^{m}_{\mathbf{M}}$ be the maximal compact open subgroup of $\mathbf{M}(\mathbb{A}_{f})$ and $K_{\mathbf{M}}$ the compact open subgroup $K\cap\mathbf{M}(\mathbb{A}_{f})$ of $\mathbf{M}(\mathbb{A}_{f})$. Let $E$ be the splitting field of $\mathbf{M}$. Since $G_{S}$ is of adjoint type, $E$ is a (Galois) CM field. Let $D_{E}$ be the absolute value of the discriminant of $E$.
@@
The discriminant of $s$ is
$\Delta(s)\coloneqq[K^{m}_{\mathbf{M}}:K_{\mathbf{M}}]D_{E}.$
We want to estimate the heights and degrees of special and pre-special points in terms of $\Delta(s).$ Fortunately, recent progress allows us to have the following proposition.
@@
Where the constants $C_{1},C_{2},\dots$ are depend on $G_{S},X,F,\mathcal{F}_{S}$ and the realization.
@@
Over $\overline{\mathbb{Q}}$, any commutative connected group $G$ can be written as a product of a semi-abelian variety with affine space $G^{\prime}\times\mathbb{G}_{a}^{m}$ (see [NW14, Prop. 5.1.12]). We will need height bounds on the semi-abelian factor, so suppose $G$ is a semi-abelian variety over a number field $L$ with toric part $T$ and abelian quotient $\pi:G\to A$. Let $g$ be the dimension of $A$. Let $K$ be any number field. Note that the Weil height on a semi-abelian variety might be negative. However, the canonical height $\widehat{h}_{L}$ on $G$ defined by Kühne [Kü20, Sec. 3], with respect to a $T$-effective line bundle $(M,\varrho)$ on a $T$-equivariant compactification $\overline{T}$ and an ample line bundle $N$ on $A$, remedies this issue. As for abelian varieties, the zero set of $\widehat{h}_{L}$ coincides with the torsion points of $G$. Let
$\eta:=\eta(G,L)=\operatorname{inf}\widehat{h}_{L}(P),$
where the infimum is taken over all non-torsion $P$ in $G(K)$, and we write
$\omega:=\omega(G,K)$
for the cardinality of the torsion group of $G(K)$. Note that $\eta$ and $\omega$ depend on the embedding of $G$ in a projective space. Suppressing this dependence in our notation, we can prove the following theorem using an argument similar to the proof of [Mas88, Theorem $\mathbb{G}_{m},A$].
@@
Suppose $P_{1},\ldots,P_{n}$ on $G(K)$ have canonical heights at most $q\geq\eta$. Then the relation group of $P_{1},\ldots,P_{n}$ generated by
$\boldsymbol{m}:m_{1}P_{1}+\cdots m_{n}P_{n}=0_{G}$
@@
Proof.
We refer to [Kü20, Lem. 8] for the definition and properties of the canonical height on a semi-abelian variety. By definition,
$\widehat{h}_{L}=\widehat{h}_{G(M,\varrho)}+\widehat{h}_{\pi^{*}N},$
where $\widehat{h}_{G(M,\varrho)}$ is linear and $\widehat{h}_{\pi^{*}N}$ is quadratic. We define a convex distance function $f$ on $\mathbb{Z}^{n}$ by
$f(\boldsymbol{m})=\widehat{h}_{G(M,\varrho)}(\boldsymbol{m})+\widehat{h}_{\pi^{*}N}(\boldsymbol{m})^{1/2}.$
Let $\Gamma=\Gamma(f)$ be the set consisting of all $\boldsymbol{m}$ such that $m_{1}P_{1}+\ldots+m_{n}P_{n}$ is a torsion point of $G(K)$, and take $\Gamma_{0}$ as the relation group of $P_{1},\ldots,P_{n}$. We can choose $E=q,\varepsilon=\eta^{1/2}$, and the theorem follows immediately from [Mas88, Prop.].
General semi-abelian subvarieties of $G$ are generated by relation groups where the $m_{i}\in\operatorname{End}(G)$. We accommodate for $\operatorname{End}(G)\neq\mathbb{Z}$ in the following way. Let $\{1,\tau_{1},\dots,\tau_{k}\}\in\operatorname{End}(G)$ be a set of generators. Then, we can work in $G^{(k+1)n}$ with $P_{i},\tau_{1}P_{i},\cdots\tau_{k}P_{i}$ for $1\leq i\leq n$. Due to the dominance of a height associated to an ample divisor (see [Lan83, Chap. 4, Prop. 5.4] or [Poo01, Lemma 6]),
(3.0.1) $\widehat{h}_{L}(\tau_{i}(P))\leq C^{i}\widehat{h}_{L}(P)$
where each $C^{i}$ is an absolute constant depending on $\tau_{i}$ and the embedding of $G$. Let
@@
where $\overline{C}$ is taken as the maximal of $(C^{i})^{(2ng-1)/2}$ and $q:=\textup{max}_{1\leq i\leq n}\{\widehat{h}_{L}(P_{i})\}$.
To estimate 3.0.2 in terms of degree of the algebraic points, we need some estimates for $\eta,\omega$, which is given for tori and abelian varieties in [Mas88]. Let $D=[K:\mathbb{Q}]$ and $\mathcal{L}=\log(D+2)$ we have:
@@
To bound the height of rational points on a suitable definable set in Section 4, we need the following upper bound of the norm of the generating set for the linear relations satisfied by $n$ fixed $V$-images of special points in terms of their discriminants.
@@
Let $(s_{1},x_{1}),\cdots,(s_{n},x_{n})\in S\times G\times\mathbb{G}_{a}^{m}$ be points of $V$ with $s_{i}$ special and with discriminants $\Delta(s_{i})$. Define the complexity of $s=(s_{1},\dots,s_{n})$ by
@@
Then there are constants $C,C^{\prime},c^{\prime}$ depending on $S,G,V,n$ such that for $\boldsymbol{\Delta}\geq C$, there is a generating set for the linear relations satisfied by the $x_{i}\in G$ such that
@@
Since $V$ is a correspondence in $S\times G$, $(s,x)\in V$ is an algebraic point, using standard properties of heights we have that
@@
and since the projection of $V$ to $S$ is finite and both projections are dominant on each factor, we have
@@
By [Kü20, Lem. 8(a)], after fixing the embedding corresponding to the $T$-effective and ample line bundles on $G$, the differences of $\widehat{h}_{L}(P)$ and $h(P)$ are bounded globally on $G(\overline{\mathbb{Q}})$ by an absolute constant $\delta$. Therefore for $h:=\textup{max}\{h(x_{1}),\cdots,h(x_{n})\}$ sufficiently large, we have
@@
By Proposition 3.2(a) $\eta$ and $q$ are bounded in terms of $\boldsymbol{\Delta}$, and the proposition follows from Corollary 3.4.
@@ 4.1. Proof of the main theorem over $\overline{\mathbb{Q}}$
First, we prove a criterion for inclusion of complex algebraic varieties defined over $\overline{\mathbb{Q}}$ and a finiteness result for pre-special subvarieties.
@@
Let $d=\dim W$. We can find a generic point $P_{1}\times P_{2}\times\cdots\times P_{m+1}\in W^{m+1}$ such that $\operatorname{tr.deg.}_{\mathbb{Q}}\mathbb{Q}(P_{1},\dots,P_{m+1})=d(m+1)$, and we claim that these $P_{i}$ satisfy the above property. The forward direction of the if and only if is clear. Now suppose that $P_{i}\times Q\in V$ for all $i$. Viewing $P_{i}$ as a set, let $\tilde{P_{i}}\subset P_{i}$ be a minimal transcendental basis for $\operatorname{tr.deg.}_{\mathbb{Q}}\mathbb{Q}(P_{i})$. Write $Q=(q_{1},\dots,q_{n})\in\mathbb{C}^{n}$ and let $Q_{i}\subset\{1,2,\dots,n\}$ be the set of indices $j$ so that $q_{j}$ is algebraic over $\mathbb{Q}(\tilde{P_{i}})$. By construction, $\lvert\tilde{P_{i}}\rvert=d$ and the $\tilde{P_{i}}$ are algebraically independent. This implies that $Q_{i}\cap Q_{j}=\varnothing$ for $i\neq j$ and so $\sum_{i=1}^{m+1}\lvert Q_{i}\rvert\leq m$. Thus, there exists an $i$ such that $\lvert Q_{i}\rvert=0$ and for this $i$ we have
@@
Let $X=\overline{\{P_{i}\}}^{\mathrm{Zar}}$ and $Y=\overline{\{Q\}}^{\mathrm{Zar}}$ be the $\overline{\mathbb{Q}}$-Zariski closure of $P_{i}$ and $Q$. Then $\dim X=\operatorname{tr.deg.}_{\mathbb{Q}}\mathbb{Q}(P_{i})=d$ and $\dim Y=\operatorname{tr.deg.}_{\mathbb{Q}}\mathbb{Q}(Q)$. Note that the Zariski closure of $P_{i}\times Q$ has dimension $\operatorname{tr.deg.}_{\mathbb{Q}}\mathbb{Q}(P_{i},Q)=\dim X+\dim Y$, showing that the $\overline{\mathbb{Q}}$-Zariski closure of $P_{i}\times Q$ is $X\times Y$. But $X\times Y\subset V$ by definition and $X=W$ showing inclusion in the other direction. ∎
@@
Suppose $V\subset S\times G$ is an algebraic subvariety of the product of a Shimura variety with a connected commutative algebraic group. There exists a finite set $\Sigma$ of sub-Shimura datum $(H,X_{H})$ of $(G_{S},X)$ and splittings $(H^{\operatorname{ad}},X_{H}^{\operatorname{ad}})=(H_{1},X_{1})\times(H_{2},X_{2})$ such that if $W\subset V$ is exemplary, then there exist one such splitting in $\Sigma$ such that $\pi_{S}(W)$ is the image of $X_{1}\times\{x_{2}\}$, for some $x_{2}\in X_{2}$, under the uniformization map of $S$.
Moreover, if $\mathcal{V}\to W$ is a definable family of $V\subset S\times G$, then the result still holds.
@@
The result for a fixed $V\subset S\times G$ is proven by [PT22, Prop. 3.4] and [DR18, Prop. 6.10]. The version for family is given by the uniform Ax–Schanuel theorem, which is implied by the Ax–Schanuel theorem. This is proven in [ES23, Prop. 2.20]. ∎
@@
Let $S$ be a connected Shimura variety and let $G$ be a connected commutative algebraic group. Suppose $V\subset S\times G$ is a proper irreducible subvariety that is finite over $S$. Moreover, suppose that $G$ and $V$ are defined over $\overline{\mathbb{Q}}$. Then, there are only finitely many exemplary components in $V$.
@@
Let $\pi_{S}\colon V\to S$ and $\pi_{G}\colon V\to G$ be the projections onto the two factors. Let $W\subset V$ be an exemplary component and let $S^{\prime}\coloneqq\pi_{S}(W)$ be the special subvariety of $S$ which $W$ maps onto, and let $G^{\prime}\supset\pi_{G}(W)$ be the smallest special subvariety of $G$ containing $\pi_{G}(W)$. $S^{\prime}$ is a special subvariety of $S$ and hence there is a sub-Shimura datum $(H,X_{H})$ of $(G_{S},X)$, a decomposition $(H^{\operatorname{ad}},X_{H}^{\operatorname{ad}})=(H_{1},X_{1})\times(H_{2},X_{2})$, and a point $y_{2}\in X_{2}$ so that $S^{\prime}$ is the image of $X_{1}\times\{y_{2}\}$. By Lemma 4.2, there are only finitely many choices of $X_{H}$ and $X_{1}$. Thus, to show finiteness of exemplary components, it suffices to show that for a fixed $X_{1}$ and fixed dimension of $G^{\prime}$, there are only finitely many points $y_{2}\in X_{2}$ arising from exemplary subvarieties.
Let $\tilde{q}\colon X_{1}\times X_{2}\to X\to S$ denote the uniformization map of $S$ restricted to $X_{1}\times X_{2}$ and let $q\colon F_{1}\times F_{2}\to S$ denote the restriction of $\tilde{q}$ to a fundamental domain $F_{1}\times F_{2}$ of $X_{1}\times X_{2}$. Let $\exp\colon\mathbb{C}^{g}\to G$ denote the uniformization map of the algebraic group and let $e\colon F_{G}\to G$ denote the restriction of that map to a fundamental domain. Using Lemma 4.1, let $x_{1},\dots,x_{r}\in F_{1}$ denote a set of points to determine if $q(X_{1}\times\{x_{2}\})\times\exp(z)\subset V$. Let $\xi_{1},\dots,\xi_{g}$ denote the coordinates in $G$ and let $\tau_{1},\dots,\tau_{k}\in\operatorname{End}(G)$ be a $\mathbb{Z}$-basis for $\operatorname{End}(G)$. Take a generating set of all equations of the form
@@
that all the points of $G^{\prime}$ satisfy, and let $G_{0}$ be the identity component of the algebraic subgroup of $G$ defined by these equations. We can extend each map $\tau_{i}\colon G\to G$ to an endomorphism of its covering space $\tilde{\tau_{i}}\colon\mathbb{C}^{g}\to\mathbb{C}^{g}$ satisfying $\tilde{\tau_{i}}(0)=0$. Suppose that $G_{0}$ is cut out by $\ell$ such equations. Let
$Y=$$\left\{(y,z_{1},\dots,z_{g},m_{111},\dots,m_{gk\ell},b_{1},\dots,b_{\ell})\in F_{2}\times F_{G}\times\mathbb{R}^{gk\ell}\times\mathbb{R}^{\ell}:\right.$
$\left.\forall 1\leq i\leq r,(q(x_{i},y),e(\boldsymbol{z}))\in V\text{ and }\forall 1\leq l\leq\ell,\sum_{i,j}m_{ijl}\tilde{\tau_{j}}(z_{i})=b_{l}\right\}$
@@
and set $Z$ to be the projection of $Y$ to $F_{2}\times\mathbb{R}^{gk\ell}\times\mathbb{R}^{\ell}$. Both $Y$ and $Z$ are definable sets. The set $Y$ parametrizes points $y\in F_{2}$ so that a $V$-image of $X_{1}\times\{y\}$ lies within a proper special subvariety cut out by the $\boldsymbol{m}$ of $G$ (but not all choices of $\boldsymbol{m}$ and $\boldsymbol{b}$ correspond to algebraic subvarieties).
@@
Suppose for the sake of contradiction that there were infinitely many $\overline{\mathbb{Q}}$-algebraic exemplary subvarieties $W^{\prime}$ with fibers over $X_{1}\times X_{2}$, and with dimension of the smallest special subvariety of $G$ containing $\pi_{G}(W^{\prime})$ equal to a fixed dimension $d=\dim G^{\prime}$. Each one gives a $\overline{\mathbb{Q}}$-point $(y^{\prime},\boldsymbol{m},\boldsymbol{b})\in Z$, and the $\operatorname{Gal}(\overline{\mathbb{Q}}/K)$-orbits of $W^{\prime}$ also lie in $Z$ for $K$ the defining field of $S,G,\operatorname{End}(G)$. Over $\overline{\mathbb{Q}}$, any commutative connected group $G$ can be written as a product of a semi-abelian variety with affine space $G^{\prime}\times\mathbb{G}_{a}^{n}$ (see [NW14, Prop. 5.1.12]). The tuple $\boldsymbol{m}$ consists integers and for semi-abelian varieties, we may take the fundamental domain so that the real part of each $z_{i}$ is in the interval $[0,1]$. Thus, each $b_{l}$ corresponding to an equation on the semi-abelian variety is an integer bounded by $\sum\lvert m_{ijl}\rvert$, which by Proposition 3.5, is bounded by $C\lvert\Delta(y^{\prime})\rvert^{c}$. Special subvarieties of $\mathbb{G}_{a}^{n}$ are given by the equations $\sum m_{ij}\tilde{\tau_{j}}(z_{i})=0$, so $b_{l}=0$ for those equations. Thus, taking the Galois orbit of $W^{\prime}$ gives a point with different $y^{\prime}$ as well as different $b_{l}$, but the same $\boldsymbol{m}$. View $Z$ is fibered over the $\boldsymbol{m}$ variable and let $\Sigma\subset Z_{\boldsymbol{m}}$ be the set of points arising from exemplary subvarieties $W^{\prime}$ as well as their Galois conjugates. By Proposition 3.2, there are at least $C^{\prime}\lvert\Delta(y^{\prime})\rvert^{c^{\prime}}$ points of height less than $C\lvert\Delta(y^{\prime})\rvert^{c}$. Then Theorem 4.3 gives the existence of a set $R\subset Z_{\boldsymbol{m}}$ of positive dimension whose projection to $F_{2}$ is connected semialgebraic and whose projection to $\mathbb{R}^{\ell}$ is non-constant.
@@
Let $\overline{\exp}\colon\mathbb{C}^{g}/\tilde{G_{0}}\to G/G_{0}$ be the exponentiation map of $G/G_{0}$. Let $F_{G_{0}}$ denote the image of the fundamental domain $F_{G}$ under the quotient map, which will serve as a fundamental domain for $\overline{\exp}$. First, take the preimage of $R\subset Z_{\boldsymbol{m}}$ under the projection map $Y\to Z$ and then project the preimage under the map
$Y_{\boldsymbol{m}}\to F_{2}\times F_{G_{0}}\times\mathbb{R}^{\ell}\to F_{2}\times F_{G_{0}}.$